# Measurement to Meaning: A Validity-Centered Framework for AI Evaluation

### v1

@@ 1 Introduction
Clearly, asserting Claim 2 requires a much greater inferential leap from the observed evidence (IMO accuracy) than Claim 1. If we claim that good performance on IMO problems demonstrates competence in solving linear algebra questions from a textbook, the justification may be reasonable: IMO problems often involve advanced undergraduate-level techniques, including linear algebra, so proficiency in them provides reasonable evidence for the claim. However, if we claim that the system has reached human-level reasoning, the justification is much weaker. Solving IMO problems primarily requires mathematical problem-solving, but human reasoning encompasses a broader spectrum, including common sense, adaptability, and metacognition, which IMO performance says little about. This difference highlights that we must scrutinize an evaluation and measurement in the context of the claim we wish to support. Thus, we consider validity. Validity refers to the degree to which evidence and theory support the interpretations of test scores for proposed uses of tests (Association et al., 2014).
Historically, measurement theory studies how latent psychological concepts are quantitatively assessed, ensuring that the chosen measures accurately capture the intended constructs while maintaining validity and reliability. Validity is not inherently a property of the measurement itself—it also depends on the context of the evaluation it enables, the claims supported by that evaluation, and the potential real-world consequences of those claims (Messick, 1998, 1995; Shepard, 1993; Lissitz and Samuelsen, 2007; Cronbach and Meehl, 1955). While this view of validity is not universal (c.f.,  Borsboom et al. (2004)), it is widely held and endorsed by several professional organizations focused on measurement in educational and psychological testing Association et al. (2014). We discuss validity and its treatment in other scientific disciplines in Appendix B.
@@
    Carefully consider the object of the claim. Is it a construct—an abstract object that cannot be measured directly, like ‘mathematical reasoning’? Or is it a criterion—a directly measurable object, such as “accuracy on linear algebra questions from a textbook”?
@@
Ensuring validity primarily requires five forms of validity from psychometrics (Thorndike, 1949; Cronbach and Meehl, 1955; Messick, 1995; Association et al., 2014; Lissitz and Samuelsen, 2007), outlined in Table 2. Additionally, Section 3 and Appendix B Table LABEL:tab:validity enumerates tools to investigate and establish each form. While other forms of validity exist and are relevant [^1] Other forms of validity include, for example, face validity. Additional forms of validity are given in (Lim, 2024; Hughes, 2018; Borsboom et al., 2004; Association et al., 2014; Hunsley and Meyer, 2003; Reichardt, 2002; Bronfenbrenner, 1977; Neisser, 1976)., we identify this set as most relevant for current AI measurement validity gaps in Section 2. As part of our practically guided framework, we also differentiate between types of claims, where some forms of validity are trivially satisfied and only a subset needs to be focused on (Lissitz and Samuelsen, 2007).
@@ 2 Validity Gaps in Current AI Evaluations and Related Work
Initially, evaluations emphasized held-out test data drawn from the same distribution as training data, which is often called the i.i.d. setting, supporting content validity, where the test environment closely matched the training one. As pretraining became the norm, evaluation shifted to fine-tuning large models like those trained on ImageNet (Deng et al., 2009; Russakovsky et al., 2014; Kornblith et al., 2018; Recht et al., 2019) and measuring downstream performance. This introduced criterion validity, where success on downstream tasks was taken as evidence that pretraining had captured useful representations (Mikolov et al., 2013; Brown et al., 2020; Radford et al., 2021).
As concerns about spurious correlations driven decision-making, distribution shifts (Bai et al., 2025; Salaudeen and Koyejo, 2024; Lopez-Paz et al., 2016; Xiao et al., 2020; Arjovsky et al., 2019; Rosenfeld et al., 2020; Koh et al., 2020; Gulrajani and Lopez-Paz, 2020), and causal representations (Schölkopf et al., 2021; Gichoya et al., 2022) rose, so too did other forms of validity. Recent work has increasingly emphasized external validity, whether evaluation results generalize beyond the training distribution (Salaudeen and Hardt, 2024), consequential validity, whether model deployment leads to desirable outcomes (Wang and Russakovsky, 2023), and construct validity, whether an evaluation actually captures the concept we think it does (Bell et al., 2024; Salaudeen et al., 2025).
Still, evaluation remains largely benchmark-driven, often in service of leaderboard-based progress (Hardt and Recht, 2021; Orr and Kang, 2024). This kind of evaluation is not without merit: when optimizers, architectures, or training procedures improve benchmark performance across multiple tasks, they also tend to improve performance on another new and real-world task, a sort of criterion validity (Blum and Hardt, 2015; Kornblith et al., 2018; Salaudeen and Hardt, 2024). Moreover, these shared benchmarks have helped align academia, industry, and other stakeholders on a criterion to measure progress (Donoho, 2023; Recht, 2024), for instance, ImageNet accuracy (Russakovsky et al., 2014). However, benchmark performance does not always translate to reliable real-world performance or trustworthy decision making (Hardy et al., 2024).
The rise of foundation models, which can operate across diverse tasks out of the box, further complicates this issue. Traditional evaluation methods increasingly fail to capture real-world AI behaviors that require investigating abstract capabilities like intelligence and reasoning to predict broad and diverse downstream utility (Wu et al., 2023; Wan et al., 2024; Mirzadeh et al., 2024).
@@
Important prior work has demonstrated the need for validity frameworks (Jacobs and Wallach, 2021; Saxon et al., 2024; Subramonian et al., 2023; Xiao et al., 2023; Blodgett et al., 2021; Coston et al., 2023; Xiao et al., 2024; Reuel et al., 2024). However, much of it has focused on validity in the context of the limitations of measurements, focusing on conditions for perfect measurements of nicely defined concepts, which is far from practice. METRICEVAL (Xiao et al., 2023) raises validity concerns stemming from vague benchmark articulation and repurposed datasets in 16 natural language generation metrics. Liu et al. (2024) further challenged benchmarks’ ability to measure intended constructs, proposing the Evidence-Centered Benchmark Design (ECBD) framework to ensure rigorous metric selection. However, these works focus on designing new measurement instruments for evaluations. While developing better measurement instruments for better evaluations is also important, and our framework also applies to this task, we find it of practical value to understand what claims can be made from existing evaluations and evidence, given the intractability of creating tailored evaluations for each claim, and the already unwieldy amount of existing benchmarks (Copilot, 2024).
@@ 3 Risks to Validity and Operationalizable Strategies for Mitigation
Risks to content validity include coverage deficiency, where important aspects of the construct are missing, and construct irrelevance, where extraneous factors influence scores (Association et al., 2014; Messick, 1995). Imbalanced content can lead to assessments overemphasizing certain skills while neglecting others. These issues can be examined through expert review, adversarial scrutiny, and synthetic data generation, with supporting evidence from explicit content mapping and coverage analysis.
Risks to external validity include sample bias, where the test is validated on a narrow or unrepresentative population (Henrich et al., 2010), and unrealistic testing conditions, which may not reflect real-world scenarios (Donald T. Campbell, 1963). Temporal variability and interaction effects can also distort results if performance shifts over time or due to specific environmental factors (Andonov et al., 2023). These issues can be investigated through stress testing, A/B testing, transfer testing, and population-stratified assessments, with evidence from performance comparisons across different conditions and sensitivity analyses.
@@
Risks to construct validity can come from structural, convergent, and discriminant validity risks. Structural validity is compromised by poor factor structure, where test items fail to group in expected ways (Clark and Watson, 1995; Elhami Athar, 2023), and complex measurement range, where constructs are not well captured across different levels of ability (Messick, 1995). Convergent validity can suffer from high measurement error (Cheung et al., 2024), which reduces reliability, while discriminant validity can be compromised by construct overlap, where different abilities are not clearly distinguished (Shaffer et al., 2016). These risks can be investigated using hypothesis testing, factor modeling, and benchmark suites, with supporting evidence from item-test correlations and demonstrated non-significant overlap with unrelated constructs.
Risks to consequential validity include bias and fairness issues, where results systematically disadvantage certain groups (Meredith, 1993; Messick, 1995; Randall, 2023). While bias and fairness can themselves be constructs of interest, they are also important to consider in any measurement. Further, unintended incentives can distort behavior if assessment criteria encourage gaming rather than genuine learning (Nichols and Berliner, 2007). Policy consequences may emerge if flawed assessments influence high-stakes decisions. These risks can be assessed through anticipatory ethics methods (Umbrello et al., 2023), societal impact audits, and ethical stress testing, with evidence from stakeholder feedback, improvements in fairness and reliability, and documented real-world impacts.
@@ 5 Application of our Framework
5.1 Scenario 1: The object of the claim is a criterion and is measured as evidence
@@
This setting is commensurate with traditional AI benchmarking practices. Many AI benchmarks have focused on these forms of generalization, including classical generalization (Vapnik and Chervonenkis, 1971) and out-of-distribution generalization (Shimodaira, 2000). By ensuring strong content and external validity, such benchmarks provide a solid foundation for validating claims for known and directly measurable criteria.
5.2 Scenarios 2-3: The object of the claim is a criterion, but a different object is measured as evidence
@@
5.3 Scenario 4. The object of the claim is a construct
@@
  • Establish Construct Validity: Performance on GPQA aligns with success in structured question-answering tasks, suggesting some reasoning component. However, structural validity is unclear, as the test may not sufficiently capture the rank of reasoning. Convergent validity is unverified since GPQA accuracy has not been correlated with other explicit reasoning assessments, e.g., interactive human evaluation. Discriminant validity is also uncertain, as it remains unclear whether GPQA measures genuine scientific reasoning or simply domain-specific knowledge and memorization. Comparing performance to humans with access to Google is an attempt to do this, though the time constraints on answering the question may add some construct-irrelevant hardness relative to human-AI comparisons. To address these concerns, methods like factor analysis (Kim and Mueller, 1979) should be conducted to distinguish reasoning from memorization, for instance, and performance should be validated against other dedicated reasoning assessments.
@@ 6 Who Should Care?
  • Policy Makers. As benchmarks increasingly guide regulatory decisions (e.g., under the EU AI Act), our framework provides a tool to assess whether the measurements being used actually support the claims that matter, like risk, safety, or societal impact. It helps ensure policy is evidence-aligned, not just benchmark-driven.
  • Corporations. Product claims, deployment readiness, and resource allocation hinge on evaluation. Our framework helps internal teams diagnose whether current evaluations support the intended claims and identify what additional evidence is needed, reducing wasted effort and increasing confidence in deployment decisions.
@@ Appendix B Validity
Moving toward empirical rigor, predictive validity assesses a test’s ability to forecast an outcome of interest, typically a future outcome. This concept, introduced by Robert Thorndike in the mid-20th century during the rise of standardized testing, became central to fields like educational assessment, employment testing, and aptitude measurement (Thorndike, 1949). For example, the predictive validity of SAT scores for college GPA or cognitive ability tests for job performance has led to their widespread use for other outcomes (Kobrin et al., 2008). In the context of AI evaluation, one may ask “Does accuracy on IMO benchmarks predict accuracy in textbook linear algebra questions?” While predictive validity measures the correlation between a test and a future outcome, concurrent validity measures the correlation between a test and a validated standard applied at the same time under the same conditions. Predictive and concurrent validity make up criterion validity (Association et al., 2014).
@@ C.1 Vision
In the 1990s, AI benchmarks became more structured and began incorporating more applied tasks. A notable example is MNIST (Lecun et al., 1998) for handwritten digit classification, which provided a standardized way to evaluate machine learning models. This trend continued into the early 2000s, with datasets such as UIUC Cars (Agarwal and Roth, 2002) for vehicle detection and Caltech-101 (2003) (Fei-Fei et al., 2005) for object recognition. While these benchmarks remained narrow in scope, they represented a step toward evaluating AI on more applied tasks, bridging the gap between theoretical research and practical applications. However, evaluations were still primarily designed for well-defined technical interests, with conclusions remaining local—focused on determining which techniques were most effective for the specific task being evaluated. During this period, researchers also became increasingly aware of content validity, recognizing that different datasets captured different aspects of classification tasks, which in turn influenced dataset design and evaluation methodologies (Ponce et al., 2006; Torralba and Efros, 2011).
@@
By the mid-2000s, large-scale benchmarks such as PASCAL VOC (2007) (Everingham et al., 2010) introduced greater complexity, expanding evaluation beyond simple classification tasks. Later, in the late 2000s, CIFAR-10 and CIFAR-100 (Krizhevsky and Hinton, 2009) further pushed the field toward standardized comparisons in object recognition. During this period, criterion validity also gained prominence, as benchmark results were increasingly used to compare models in ways that suggested performance rankings carried external significance. However, construct validity remained largely unexplored—models were evaluated based on their outputs rather than on the reasoning processes behind their decisions. As a result, while evaluations became more sophisticated, they remained focused on performance metrics rather than deeper insights into model behavior. By this stage, the focus of AI evaluation began shifting from isolated dataset-specific improvements to broader claims about model robustness and transferability across different domains.
@@
The 2010s marked a turning point with the ImageNet revolution. The introduction of ImageNet (Deng et al., 2009) and the ILSVRC (Russakovsky et al., 2014) competition (2010) provided large-scale, diverse, and complex benchmarks that dramatically reshaped AI research. During the early 2010s, the focus remained on improving accuracy in image classification and object detection. However, by the mid-2010s, AI evaluation expanded beyond leaderboards to real-world applications, particularly in medical imaging and autonomous driving. Researchers increasingly recognized the importance of content validity and external validity, leading to the widespread practice of testing models across multiple datasets to assess robustness.
@@
As benchmark results gained influence, criterion validity became central—accuracy on ImageNet was frequently treated as a proxy for predicting downstream AI capabilities in vision. However, construct validity remained largely unaddressed in the early years. By the mid-2010s, early concerns emerged as researchers identified shortcut learning, adversarial vulnerabilities, and spurious correlations, leading to growing interest in understanding how models made decisions beyond raw accuracy. The rise of segmentation (COCO (Lin et al., 2014), ADE20K (Zhou et al., 2016)) and video analysis benchmarks (Kinetics, AVA) reflected an effort to capture more complex real-world tasks, but fundamental concerns about model robustness and bias persisted.

### v3

@@ 1 Introduction
Clearly, asserting Claim 2 requires a much greater inferential leap from the observed evidence (IMO accuracy) than Claim 1. If we claim that good performance on IMO problems demonstrates competence in solving linear algebra questions from a textbook, the justification may be reasonable: IMO problems often involve advanced undergraduate-level techniques, including linear algebra, so proficiency in them provides reasonable evidence for the claim. However, if we claim that the system has reached human-level reasoning, the justification is much weaker. Solving IMO problems primarily requires mathematical problem-solving, but human reasoning encompasses a broader spectrum, including common sense, adaptability, and metacognition, which IMO performance says little about. This difference highlights that we must scrutinize an evaluation and measurement in the context of the claim we wish to support. Thus, we consider validity. Validity refers to the degree to which evidence and theory support the interpretations of test scores for proposed uses of tests (American Educational Research Association et al., 2014).
Historically, measurement theory studies how latent psychological concepts are quantitatively assessed, ensuring that the chosen measures accurately capture the intended constructs while maintaining validity and reliability. Validity is not inherently a property of the measurement itself—it also depends on the context of the evaluation it enables, the claims supported by that evaluation, and the potential real-world consequences of those claims (Messick, 1998, 1995; Shepard, 1993; Lissitz and Samuelsen, 2007; Cronbach and Meehl, 1955). While this view of validity is not universal (c.f.,  Borsboom et al. (2004)), it is widely held and endorsed by several professional organizations focused on measurement in educational and psychological testing American Educational Research Association et al. (2014). We discuss validity and its treatment in other scientific disciplines in Appendix B.
@@
    Carefully consider the object of the claim. Is it a construct—an abstract object that cannot be measured directly, like ‘mathematical reasoning’? Or is it a criterion—a directly measurable object, such as “accuracy on linear algebra questions from a specific textbook”?
@@
Ensuring validity primarily requires five forms of validity from psychometrics (Electroencephalographs, 1952; Cronbach and Meehl, 1955; Messick, 1995; American Educational Research Association et al., 2014; Lissitz and Samuelsen, 2007), outlined in Table 2. Additionally, Section 3 and Appendix B Table LABEL:tab:validity enumerates tools to investigate and establish each form. While other forms of validity exist and are relevant [^1] Other forms of validity include, for example, face validity. Additional forms of validity are given in (Lim, 2024; Hughes, 2018; Borsboom et al., 2004; American Educational Research Association et al., 2014; Hunsley and Meyer, 2003; Reichardt, 2002; Bronfenbrenner, 1977; Neisser, 1976)., we identify this set as most relevant for current AI measurement validity gaps in Section 2. As part of our practically guided framework, we also differentiate between types of claims, where some forms of validity are trivially satisfied and only a subset needs to be focused on (Lissitz and Samuelsen, 2007).
@@ 2 Validity Gaps in Current AI Evaluations and Related Work
Initially, evaluations emphasized held-out test data drawn from the same distribution as training data, which is often called the i.i.d. setting, supporting content validity, where the test environment closely matched the training one. As pretraining became the norm, evaluation shifted to fine-tuning large models like those trained on ImageNet (Deng et al., 2009; Russakovsky et al., 2015; Kornblith et al., 2018; Recht et al., 2019) and measuring downstream performance. This introduced criterion validity, where success on downstream tasks was taken as evidence that pretraining had captured useful representations (Mikolov et al., 2013; Brown et al., 2020; Radford et al., 2021).
As concerns about spurious correlations driven decision-making, distribution shifts (Bai et al., 2025; Salaudeen and Koyejo, 2024; Lopez-Paz et al., 2017; Xiao et al., 2020; Arjovsky et al., 2019; Rosenfeld et al., 2020; Marklund et al., 2020; Gulrajani and Lopez-Paz, 2020), and causal representations (Schölkopf et al., 2021; Gichoya et al., 2022) rose, so too did other forms of validity. Recent work has increasingly emphasized external validity, whether evaluation results generalize beyond the training distribution (Salaudeen and Hardt, 2024), consequential validity, whether model deployment leads to desirable outcomes (Wang and Russakovsky, 2023), and construct validity, whether an evaluation actually captures the concept we think it does (Bell et al., 2024; Salaudeen et al., 2025).
Still, evaluation remains largely benchmark-driven, often in service of leaderboard-based progress (Hardt and Recht, 2021; Orr and Kang, 2024). This kind of evaluation is not without merit: when optimizers, architectures, or training procedures improve benchmark performance across multiple tasks, they also tend to improve performance on another new and real-world task, a sort of criterion validity (Blum and Hardt, 2015; Kornblith et al., 2018; Salaudeen and Hardt, 2024). Moreover, these shared benchmarks have helped align academia, industry, and other stakeholders on a criterion to measure progress (Donoho, 2024; Recht, 2024), for instance, ImageNet accuracy (Russakovsky et al., 2015). However, benchmark performance does not always translate to reliable real-world performance or trustworthy decision making (Hardy et al., 2024).
The rise of foundation models, which can operate across diverse tasks out of the box, further complicates this issue. Traditional evaluation methods increasingly fail to capture real-world AI behaviors that require investigating abstract capabilities like intelligence and reasoning to predict broad and diverse downstream utility (Wu et al., 2024; Wan et al., 2024; Mirzadeh et al., 2024).
@@
Important prior work has demonstrated the need for validity frameworks (Jacobs and Wallach, 2019; Saxon et al., 2024; Subramonian et al., 2023; Xiao et al., 2023; Blodgett et al., 2021; Coston et al., 2022; Liu et al., 2025; Reuel et al., 2024). However, much of it has focused on validity in the context of the limitations of measurements, focusing on conditions for perfect measurements of nicely defined concepts, which is far from practice. METRICEVAL (Xiao et al., 2023) raises validity concerns stemming from vague benchmark articulation and repurposed datasets in 16 natural language generation metrics. Liu et al. (2024) further challenged benchmarks’ ability to measure intended constructs, proposing the Evidence-Centered Benchmark Design (ECBD) framework to ensure rigorous metric selection. However, these works focus on designing new measurement instruments for evaluations. While developing better measurement instruments for better evaluations is also important, and our framework also applies to this task, we find it of practical value to understand what claims can be made from existing evaluations and evidence, given the intractability of creating tailored evaluations for each claim, and the already unwieldy amount of existing benchmarks (Paper Copilot, 2024).
@@ 3 Risks to Validity and Operationalizable Strategies for Mitigation
Risks to content validity include coverage deficiency, where important aspects of the construct are missing, and construct irrelevance, where extraneous factors influence scores (American Educational Research Association et al., 2014; Messick, 1995). Imbalanced content can lead to assessments overemphasizing certain skills while neglecting others. These issues can be examined through expert review, adversarial scrutiny, and synthetic data generation, with supporting evidence from explicit content mapping and coverage analysis.
Risks to external validity include sample bias, where the test is validated on a narrow or unrepresentative population (Henrich et al., 2010), and unrealistic testing conditions, which may not reflect real-world scenarios (Ruse, 2019). Temporal variability and interaction effects can also distort results if performance shifts over time or due to specific environmental factors (Andonov et al., 2023). These issues can be investigated through stress testing, A/B testing, transfer testing, and population-stratified assessments, with evidence from performance comparisons across different conditions and sensitivity analyses.
@@
Risks to construct validity can come from structural, convergent, and discriminant validity risks. Structural validity is compromised by poor factor structure, where test items fail to group in expected ways (Clark and Watson, 1995; Athar, 2023), and complex measurement range, where constructs are not well captured across different levels of ability (Messick, 1995). Convergent validity can suffer from high measurement error (Cheung et al., 2024), which reduces reliability, while discriminant validity can be compromised by construct overlap, where different abilities are not clearly distinguished (Shaffer et al., 2016). These risks can be investigated using hypothesis testing, factor modeling, and benchmark suites, with supporting evidence from item-test correlations and demonstrated non-significant overlap with unrelated constructs.
Risks to consequential validity include bias and fairness issues, where results systematically disadvantage certain groups (Meredith, 1993; Messick, 1995; Randall, 2023). While bias and fairness can themselves be constructs of interest, they are also important to consider in any measurement. Further, unintended incentives can distort behavior if assessment criteria encourage gaming rather than genuine learning (Callet, 2008). Policy consequences may emerge if flawed assessments influence high-stakes decisions. These risks can be assessed through anticipatory ethics methods (Umbrello et al., 2023), societal impact audits, and ethical stress testing, with evidence from stakeholder feedback, improvements in fairness and reliability, and documented real-world impacts.
@@ 5 Application of our Framework
5.1 Criterion-Aligned Evidence (1): The object of the claim is a criterion and is measured as evidence
@@
This setting is commensurate with traditional AI benchmarking practices. Many AI benchmarks have focused on these forms of generalization, including classical generalization (Vapnik and Chervonenkis, 2015) and out-of-distribution generalization (Shimodaira, 2000). By ensuring strong content and external validity, such benchmarks provide a solid foundation for validating claims for known and directly measurable criteria.
5.2 Criterion-Adjacent Evidence (2-3): The object of the claim is a criterion, but a different object is measured as evidence
@@
5.3 Construct-Targeted Evidence (4): The object of the claim is a construct
@@
  • Establish Construct Validity: Performance on GPQA aligns with success in structured question-answering tasks, suggesting some reasoning component. However, structural validity is unclear, as the test may not sufficiently capture the rank of reasoning. Convergent validity is unverified since GPQA accuracy has not been correlated with other explicit reasoning assessments, e.g., interactive human evaluation. Discriminant validity is also uncertain, as it remains unclear whether GPQA measures genuine scientific reasoning or simply domain-specific knowledge and memorization. Comparing performance to humans with access to Google is an attempt to do this, though the time constraints on answering the question may add some construct-irrelevant hardness relative to human-AI comparisons. To address these concerns, methods like factor analysis (Kim and Mueller, 1978) should be conducted to distinguish reasoning from memorization, for instance, and performance should be validated against other dedicated reasoning assessments.
@@ 6 Who Should Care?
  • Policy Makers. As benchmarks increasingly guide regulatory decisions (e.g., under the EU AI Act), our framework provides a tool to assess whether the measurements being used actually support the claims that matter, like risk, safety, or societal impact. It helps ensure policy is evidence-aligned.
  • Corporations. Product claims, deployment readiness, and resource allocation hinge on evaluation. Our framework helps internal teams diagnose whether current evaluations support the intended claims or identify what additional evidence is needed, reducing wasted effort and increasing confidence in deployment decisions.
@@ Appendix B Validity
Moving toward empirical rigor, predictive validity assesses a test’s ability to forecast an outcome of interest, typically a future outcome. This concept, introduced by Robert Thorndike in the mid-20th century during the rise of standardized testing, became central to fields like educational assessment, employment testing, and aptitude measurement (Electroencephalographs, 1952). For example, the predictive validity of SAT scores for college GPA or cognitive ability tests for job performance has led to their widespread use for other outcomes (Kobrin et al., 2008). In the context of AI evaluation, one may ask “Does accuracy on IMO benchmarks predict accuracy in textbook linear algebra questions?” While predictive validity measures the correlation between a test and a future outcome, concurrent validity measures the correlation between a test and a validated standard applied at the same time under the same conditions. Predictive and concurrent validity make up criterion validity (American Educational Research Association et al., 2014).
@@ C.1 Vision
In the 1990s, AI benchmarks became more structured and began incorporating more applied tasks. A notable example is MNIST (LeCun et al., 2002) for handwritten digit classification, which provided a standardized way to evaluate machine learning models. This trend continued into the early 2000s, with datasets such as UIUC Cars (Agarwal and Roth, 2002) for vehicle detection and Caltech-101 (2003) (Fei-Fei et al., 2004) for object recognition. While these benchmarks remained narrow in scope, they represented a step toward evaluating AI on more applied tasks, bridging the gap between theoretical research and practical applications. However, evaluations were still primarily designed for well-defined technical interests, with conclusions remaining local—focused on determining which techniques were most effective for the specific task being evaluated. During this period, researchers also became increasingly aware of content validity, recognizing that different datasets captured different aspects of classification tasks, which in turn influenced dataset design and evaluation methodologies (Ponce et al., 2006; Torralba and Efros, 2011).
@@
By the mid-2000s, large-scale benchmarks such as PASCAL VOC (2007) (Everingham et al., 2014) introduced greater complexity, expanding evaluation beyond simple classification tasks. Later, in the late 2000s, CIFAR-10 and CIFAR-100 (Krizhevsky and Hinton, 2009) further pushed the field toward standardized comparisons in object recognition. During this period, criterion validity also gained prominence, as benchmark results were increasingly used to compare models in ways that suggested performance rankings carried external significance. However, construct validity remained largely unexplored—models were evaluated based on their outputs rather than on the reasoning processes behind their decisions. As a result, while evaluations became more sophisticated, they remained focused on performance metrics rather than deeper insights into model behavior. By this stage, the focus of AI evaluation began shifting from isolated dataset-specific improvements to broader claims about model robustness and transferability across different domains.
@@
The 2010s marked a turning point with the ImageNet revolution. The introduction of ImageNet (Deng et al., 2009) and the ILSVRC (Russakovsky et al., 2015) competition (2010) provided large-scale, diverse, and complex benchmarks that dramatically reshaped AI research. During the early 2010s, the focus remained on improving accuracy in image classification and object detection. However, by the mid-2010s, AI evaluation expanded beyond leaderboards to real-world applications, particularly in medical imaging and autonomous driving. Researchers increasingly recognized the importance of content validity and external validity, leading to the widespread practice of testing models across multiple datasets to assess robustness.
@@
As benchmark results gained influence, criterion validity became central—accuracy on ImageNet was frequently treated as a proxy for predicting downstream AI capabilities in vision. However, construct validity remained largely unaddressed in the early years. By the mid-2010s, early concerns emerged as researchers identified shortcut learning, adversarial vulnerabilities, and spurious correlations, leading to growing interest in understanding how models made decisions beyond raw accuracy. The rise of segmentation (COCO (Lin et al., 2014), ADE20K (Zhou et al., 2019)) and video analysis benchmarks (Kinetics, AVA) reflected an effort to capture more complex real-world tasks, but fundamental concerns about model robustness and bias persisted.